# Non-linear Welfare-Aware Strategic Learning

### v2

@@ 3.2 Welfare of different parties
For agent safety, we define it as:
@@ 4.1 Beyond the linear setting
Bechavod et al. (2022) considered linear settings when both the labeling function $h$ and decision policy $f$ are linear and revealed that decision-maker welfare ($\text{DW}(f)$) and social welfare ($\text{IMP}(f)$) can attain the optimum simultaneously when the decision maker deploys $f=h$ and agents best respond to $f$. However, under general non-linear settings, the compatibility of welfare is rarely achieved and we present the necessary and sufficient conditions to guarantee the alignment of all welfare in Thm. 4.1.
@@
Suppose agents have information level $K$, then decision-maker welfare, social welfare, and agent welfare can attain maximum simultaneously regardless of the feature distribution $P_{X}$ and cost function $c(x,x^{\prime})$ if and only if $h$ is polynomial in $X$ with order smaller than or equal to $K$ and $h\in\mathcal{F}$.
Thm. 4.1 can be proved based on Prop. 3.3. Intuitively, decision-maker welfare and agent welfare are guaranteed to be maximized when $f$ is accurate, while agents are only guaranteed to gain the largest improvement under an accurate $f$ if they have a perfect understanding of $f$. When Thm. 4.1 is not satisfied, the welfare of different parties is incompatible under the non-linear settings; we use two examples to illustrate this below.
@@
Under the same setting of Example 4.2, instead of restricting $f$ to be linear, let the decision-maker deploy $f(x)=h(x)=-4x\cdot(x-1)$. Obviously, $f$ maximizes both the decision-maker welfare and agent welfare. However, it can hurt social welfare. Consider an example shown in Fig. 2 where all agents have the same feature $x=0.4$ and information level $K=1$ (violating the information level requirement of Thm. 4.1). Their cost function is $c(x,x^{\prime})=(x^{\prime}-x)^{2}$. Since $\nabla f(x)=-8x+4=0.8$, agents derive $Q^{f}_{x}(x^{\prime})=f(x)+0.8\cdot(x^{\prime}-x)=0.8x^{\prime}+0.64$. Then agent post-response feature will be $x^{*}=x+\frac{0.8}{2}=0.8$. As a result, the agent qualification becomes $h(x^{*})=h(0.8)=0.64<h(0.4)=0.96$, which is worse than the initial qualification, i.e., agent safety is not optimal by deploying $f=h$.
@@ 4.2 Compatibility analysis of each welfare pair
The decision-maker welfare and social welfare are guaranteed to attain the optimum simultaneously regardless of the feature distribution $P_{X}$ and the cost function $c(x,x^{\prime})$ if and only if Thm. 4.1 holds.
@@
While Prop. 4.5 holds for both measures (i.e., agent improvement $\text{IMP}(f)$ and agent safety $\text{SF}(f)$) of social welfare, the two in general are not compatible (as shown in Example 4.3). However, we can identify the necessary and sufficient conditions under which $\text{IMP}(f)$ and $\text{SF}(f)$ attain the maximum simultaneously in general non-linear settings.
@@
Suppose agents have information level $K$ and respond to $f$ based on Def. 3.2. Let $^{*}\in\arg\max_{f\in\mathcal{F}}\text{IMP(f)}$(f) be any policy that maximizes the agent improvement, and let $Q^{f^{*}}_{x}$ be the estimation of $f^{*}$ for agents with feature vector $x$. Then any $f^{*}$ is guaranteed to be safe (i.e., $\text{SF}(f^{*})=0$) regardless of the feature distribution $P_{X}$ and the cost function $c(x,x^{\prime})$ if and only if either of the following scenarios holds: (i) Thm. 4.1 holds; (ii) For each agent with feature vector $x$, we have: $\frac{\partial h}{\partial\widetilde{x}_{i}}\cdot\frac{\partial Q^{f^{*}}_{x}}{\partial\widetilde{x}_{i}}\geq 0$ for any $\widetilde{x}=[\widetilde{x}_{1},...,\widetilde{x}_{d}]\in\text{dom}(X)$.
It is trivial to see scenario (i) in Thm. 4.6 is a sufficient condition (it ensures the compatibility of three welfare). For scenario (ii), it implies that for any arbitrary agent with feature $x$, the $f^{*}$ will result in an estimation $Q^{f^{*}}_{x}$, under which each agent changes her features in directions to improve the qualification.
@@
Suppose agents have information level $K$ and respond to $f$ based on Def. 3.2. Denote =$\mathcal{F}^{\prime}=\{f:f\in\mathcal{F}~{}and~{}f(x)\geq h(x),\forall x\in\text{dom}(X)\}$ as the set of policies maximizing the agent welfare. Then for any $f^{*}\in\arg\max_{f\in\mathcal{F}}\text{IMP}(f)$, $f^{*}$ is guaranteed to also maximize $\text{AW}(f)$ regardless of the feature distribution $P_{X}$ and the cost function $c(x,x^{\prime})$ if and only if either of the following scenarios holds: (i) Thm. 4.1 holds; (ii) $\exists f^{\prime}\in\mathcal{F}^{\prime}$ and constant $C\in\operatorname{\mathbb{R}}^{+}$ such that $Q^{f^{\prime}}_{x}=Q^{f^{*}}_{x}+C$ for any $x\in\text{dom}(X)$.
Again, scenario (i) in Thm. 4.7 is a sufficient condition, while scenario (ii) means that the decision-maker can find a classifier that induces perfect agent welfare while simultaneously giving the agents the same information as $f^{*}$. Note that $Q^{f^{\prime}}_{x}=Q^{f^{*}}_{x}+C$ not necessarily implies $f^{\prime}=f^{*}+C$ when $\text{dom}(X)$ is a zero-measure set. For example, when $K=2$ and $\text{dom}(X)=\{0\}$, we can have $f^{*}=\text{e}^{x}-x-1$ and $f^{\prime}=\frac{x^{2}}{2}+0.5$, where $f^{*}$ belongs to the exponential function family but $f^{\prime}$ belongs to the quadratic function family.
@@ 6.1 Datasets
  • ACSIncome-CA data (Ding et al. 2021). It contains income-related data for California residents. Let label $Y$ be whether an agent’s annual income is higher than $50$K USD and let the sensitive attribute $Z$ be sex. We randomly select $20000$ samples from the data of year $2018$. We assume all agents have the same effort budget $c(x,x^{\prime})=\|x^{\prime}-x\|^{2}$.
  • German Credit data (Dua and Graff 2017). It has $1000$ samples and the label $Y$ is whether the agent’s credit risk is high or low. We adopt the data preprocessed by (Jiang and Nachum 2020) where age serves as the sensitive attribute to split agents into two groups. We assume the agents are able to improve the following features: existing account status, credit history, credit amount, saving account, present employment, installment rate, guarantors, residence. We assume all agents have the same effort budget $c(x,x^{\prime})=\|x^{\prime}-x\|^{2}$.
@@ 6.3 Comparison with Previous Algorithms
Equal Improvability (EI) $P\left(f(x^{*})\geq 0.5\mid f(x)<0.5,Z=z\right)=P\left(f(x^{*})\geq 0.5\mid f(x)<0.5\right)$
@@ Acknowledgement
Appendix A Proofs
A.1 Proof of Prop. 3.3
@@ Appendix A Proofs
1. Proof of the necessity: Since the maximization of the decision-maker welfare regardless of $P_{X}$ needs $f=h$ (otherwise, one can always take the point mass on the point that $f$ misidentifies as the new feature distribution and $f$ is no longer the one maximizing the decision-maker welfare), then $h\in\mathcal{F}$ is needed. Then the agents need to best respond to $f$ to ensure the maximization of social welfare . If there is an arbitrary agent who does not best respond to $f=h$, this means she misses the largest possible improvement and the overall maximization of social welfare is failed. Thus, according to Prop. 3.3, $h$ itself must be a polynomial of $X$ with at most order $K$ to guarantee agents to best respond.
2. Proof of sufficiency: When $h$ is a polynomial with order at most $K$ and $h\in\mathcal{F}$, the decision-maker can directly deploy $h=f$ to ensure the maximization of all welfare.
A.3 Proof of Prop. 4.5
@@
1. Proof of the necessity: Since the maximization of the decision-maker welfare needs $f=h$, then $h\in\mathcal{F}$ is needed. Then the agents need to best respond to $f$ to ensure the maximization of social welfare. If there is an arbitrary agent who does not best respond to $f=h$, this means she misses the largest possible improvement and the overall maximization of social welfare is failed. Thus, according to Prop. 3.3, $h$ itself must be a polynomial of $X$ with at most order $K$ to guarantee agents to best respond.
@@
1. Proof of the necessity: When (i) does not hold, $f^{*}$ is not equal to $h$ and the agents cannot perfectly estimate $f^{*}$ with $Q^{f^{*}}_{x}$. Therefore, if (ii) does not hold, this means there must exist a cost function $c(x,x^{\prime})$, under which we can find a region $A\subset\text{dom}(X)$ where $Q^{f^{*}}_{x}$ increases while $h$ decreases inside $A$. Thus, if we specify a cost function and an agent to satisfy the requirement that the agent can only change her features within $A$ and when $P_{X}$ is any continuous distribution, then $f^{*}$ is unsafe, and we prove that either (i) or (ii) should at least hold.
2. Proof of the sufficiency: (i) is already a sufficient condition. Then for (ii), we know for each dimension $i$, $Q^{f^{*}}_{x}$ improves at the same direction as $h$ and only the magnitude differs. This means agents will never be misguided by $Q^{f^{*}}_{x}$, thereby ensuring the safety. ∎
A.5 Proof of Thm. 4.7
@@
Appendix B Plots with error bars

### v3

@@ 3.2 Welfare of different parties
However, maximizing improvement should be based on ensuring safety. For agent safety, we define it as:
@@ 4.1 Beyond the linear setting
Bechavod et al. (2022) considered linear settings when both the labeling function $h$ and decision policy $f$ are linear and revealed that decision-maker welfare ($\text{DW}(f)$) and agent improvement ($\text{IMP}(f)$) can attain the optimum simultaneously when the decision maker deploys $f=h$ and agents best respond to $f$ with a quadratic cost. However, under general non-linear settings, the compatibility of welfare is rarely achieved and we present the necessary and sufficient conditions to guarantee the alignment of all welfare in Thm. 4.1.
@@
Suppose agents have information level $K$ and $c(x,x^{\prime})=(x-x^{\prime})^{T}Q(x-x^{\prime})$ where $Q$ is positive definite, then decision-maker welfare, social welfare, and agent welfare attain maximum simultaneously regardless of the feature distribution $P_{X}$ and $Q$ if and only if: (i) $h=\hat{w}^{T}x\in\mathcal{F}=\{x\rightarrow w^{T}x|w^{T}Rw=c^{2}\}$ where $R\in\mathbb{R}^{d\times d}$ is symmetric and positive definite and $c\in\mathbb{R}^{+}$ is a constant; (ii) $\hat{w}$ must be an eigenvector of $R^{-1}Q^{-1}$; (iii) $K\geq 1$.
Thm. 4.1 can be proved based on Prop. 3.3 together with considerations on maximizing $\text{IMP}(f)$. Intuitively, decision-maker welfare and agent welfare are guaranteed to be maximized when $f$ is accurate, while safety is only ensured under an accurate $f$ if agents have a perfect understanding of $f$. Moreover, to ensure $\text{IMP}(f)$ is maximized when $f=h$, the function class $\mathcal{F}$ must be restricted to be an ellipsoidal surface. When Thm. 4.1 is not satisfied, the welfare of different parties is incompatible under the non-linear settings; we use two examples to illustrate this below.
@@
Under the same setting of Example 4.2, instead of restricting $f$ to be linear (i.e., violating Thm. 4.1), let the decision-maker deploy $f(x)=h(x)=-4x\cdot(x-1)$. Obviously, $f$ maximizes both the decision-maker welfare and agent welfare. However, it can hurt social welfare. Consider an example shown in Fig. 2 where all agents have the same feature $x=0.4$ and information level $K=1$. Their cost function is $c(x,x^{\prime})=(x^{\prime}-x)^{2}$. Since $\nabla f(x)=-8x+4=0.8$, agents derive $Q^{f}_{x}(x^{\prime})=f(x)+0.8\cdot(x^{\prime}-x)=0.8x^{\prime}+0.64$. Then agent post-response feature will be $x^{*}=x+\frac{0.8}{2}=0.8$. As a result, the agent qualification becomes $h(x^{*})=h(0.8)=0.64<h(0.4)=0.96$, which is worse than the initial qualification, i.e., agent safety is not optimal by deploying $f=h$.
@@ 4.2 Compatibility analysis of each welfare pair
Suppose agents have information level $K$ and $c(x,x^{\prime})=(x-x^{\prime})^{T}Q(x-x^{\prime})$ where $Q$ is positive definite. The decision-maker welfare and social welfare attain the optimum simultaneously regardless of the feature distribution $P_{X}$ and $Q$ if and only if Thm. 4.1 holds.
@@
While maximum social welfare means maximizing agent improvement $\text{IMP}(f)$ conditioned on perfect agent safety $\text{SF}(f)$, the two in general are not compatible (as shown in Example 4.3). However, we can identify the necessary and sufficient conditions under which $\text{IMP}(f)$ and $\text{SF}(f)$ attain the maximum simultaneously in general non-linear settings.
@@
Suppose agents have information level $K$ and respond to $f$ based on Def. 3.2. Let $f^{*}\in\arg\max_{f\in\mathcal{F}}\text{IMP(f)}$ be any policy that maximizes the agent improvement, and let $Q^{f^{*}}_{x}$ be the estimation of $f^{*}$ for agents with feature vector $x$. Then any $f^{*}$ is guaranteed to be safe (i.e., $\text{SF}(f^{*})=0$) regardless of the feature distribution $P_{X}$ and the cost function $c(x,x^{\prime})$ if and only if: for each agent with feature vector $x$, we have: $\frac{\partial h}{\partial\widetilde{x}_{i}}\cdot\frac{\partial Q^{f^{*}}_{x}}{\partial\widetilde{x}_{i}}\geq 0$ for any $\widetilde{x}=[\widetilde{x}_{1},...,\widetilde{x}_{d}]\in\text{dom}(X)$.
This condition implies that for any arbitrary agent with feature $x$, the $f^{*}$ will result in an estimation $Q^{f^{*}}_{x}$, under which each agent changes her features in directions to improve the qualification.
@@
Suppose agents have information level $K$ and $c(x,x^{\prime})=(x-x^{\prime})^{T}Q(x-x^{\prime})$ where $Q$ is positive definite. They respond to $f$ based on Def. 3.2. Denote $\mathcal{F}^{\prime}=\{f:f\in\mathcal{F}~and~f(x)\geq h(x),\forall x\in\text{dom}(X)\}$ as the set of policies maximizing the agent welfare. Then for any $^{*}\in\arg\max_{f\in\mathcal{F}}\text{IMP}(f)+\text{SF}(f), ff^{*}$ is guaranteed to also maximize $\text{AW}(f)$ regardless of the feature distribution $P_{X}$ and $Q$ if and only if either of the following scenarios holds: (i) Thm. 4.1 holds; (ii) $\exists f^{\prime}\in\mathcal{F}^{\prime}$ and constant $C\in\operatorname{\mathbb{R}}^{+}$ such that $Q^{f^{\prime}}_{x}=Q^{f^{*}}_{x}+C$ for any $x\in\text{dom}(X)$.
Scenario (i) in Thm. 4.7 is a sufficient condition, while scenario (ii) means that the decision-maker can find a classifier that induces perfect agent welfare while simultaneously giving the agents the same information as $f^{*}$. Note that $Q^{f^{\prime}}_{x}=Q^{f^{*}}_{x}+C$ not necessarily implies $f^{\prime}=f^{*}+C$ when $\text{dom}(X)$ is a zero-measure set. For example, when $K=2$ and $\text{dom}(X)=\{0\}$, we can have $f^{*}=\text{e}^{x}-x-1$ and $f^{\prime}=\frac{x^{2}}{2}+0.5$, where $f^{*}$ belongs to the exponential function family but $f^{\prime}$ belongs to the quadratic function family.
@@ 6.1 Datasets
  • ACSIncome-CA data (Ding et al. 2021). It contains income-related data for California residents. Let label $Y$ be whether an agent’s annual income is higher than 5050K USD and let the sensitive attribute $Z$ be sex. We randomly select $20000$ samples from the data of year $2018$. We assume all agents have the same cost function $c(x,x^{\prime})=\|x^{\prime}-x\|^{2}$.
  • German Credit data (Dua and Graff 2017). It has $1000$ samples and the label $Y$ is whether the agent’s credit risk is high or low. We adopt the data preprocessed by (Jiang and Nachum 2020) where age serves as the sensitive attribute to split agents into two groups. We assume the agents are able to improve the following features: existing account status, credit history, credit amount, saving account, present employment, installment rate, guarantors, residence. We assume all agents have the same cost function $c(x,x^{\prime})=\|x^{\prime}-x\|^{2}$.
@@ 6.3 Comparison with Previous Algorithms
Equal Improvability (EI) $P\left(f(x^{*})\hskip-1.42262pt\geq\hskip-1.42262pt0.5\mid f(x)\hskip-1.42262pt<\hskip-1.42262pt0.5,Z=z\hskip-1.42262pt\right)\hskip-1.42262pt=\hskip-1.42262ptP\left(f(x^{*})\hskip-1.42262pt\geq\hskip-1.42262pt0.5\mid f(x)\hskip-1.42262pt<\hskip-1.42262pt0.5\hskip-1.42262pt\right)$
@@ Acknowledgement
Appendix A Notes
Compared with the conference version, this draft clarifies that the cost functions in the welfare analysis are quadratic, which was previously an implicit assumption. We also add $2$ missing conditions in Theorem 4.1, namely that $\mathcal{F}$ must be linear with weights on a ellipsoidal surface parameterized by $R$, and the weight vector of $h$ must be the eigenvector of $R^{-1}Q^{-1}$; otherwise, $\text{IMP}(f)$ is not guaranteed to be maximized when $f=h$. These clarifications do not affect the main claims or the proposed algorithms. In fact, they highlight that aligning welfare across all parties is even more challenging under these conditions, thereby motivating STWF even in a stronger way.
Appendix B Proofs
B.1 Proof of Prop. 3.3
@@ Appendix B Proofs
1. Proof of necessity: Since the maximization of the decision-maker welfare regardless of $P_{X}$ needs $f=h$ (otherwise, one can always take the point mass on the point that $f$ misidentifies as the new feature distribution and $f$ is no longer the one maximizing the decision-maker welfare), then $h\in\mathcal{F}$ is needed. Then the agents need to best respond to $f$ to ensure the safety. If there is an arbitrary agent who does not best respond to $f=h$, this means she misunderstands $h$ and this can result in deteriorating qualification. Thus, according to Prop. 3.3, $h$ itself must be a polynomial of $X$ with at most order $K$ to guarantee agents to best respond.
Next, consider the situation when $\mathcal{F}$ consists not only linear functions with normalized weights: (i) Firstly, if $\mathcal{F}$ is not linear, we can construct a counter-example: consider $x\in[0,1]$ in $1$-d space and $h(x)=x$. $Q=0.6$, and the population is a point mass at 0. Then if we deploy $f(x)=x^{2}$, then every agent with $x=0$ will change their features to $x=1$, but deploying $f(x)=x$ will only result in their changing to $x=0.833$. Therefore, $f=h$ does not result in largest $\text{IMP}(f)$. This example demonstrates that we cannot even relax $\mathcal{F}$ to the quadratic function class; (ii) Next, if $\mathcal{F}$ is linear and $h=\hat{w}^{T}x$, $f=w^{T}x$, we first solve the best response as follows:
Let $d:=x^{\prime}-x$. We need to derive $\max_{d\in\mathbb{R}^{d}}\ \ w^{\top}d-d^{\top}Qd$ and this gives $\nabla_{d}\big{(}w^{\top}d-d^{\top}Qd\big{)}=w-2Qd=0\Longrightarrow\ d^{*}(w)=\frac{1}{2}Q^{-1}w$. Thus, $x^{\prime}(w)=x+d^{*}(w)=x+\frac{1}{2}Q^{-1}w$. Next, $\text{IMP}(f)=\frac{1}{2}(Q^{-1}\hat{w})^{T}w$. Obviously, when $w$ is unbounded, $\text{IMP}(f)$ can be infinite. To always let $\hat{w}$ to be the maximizer of itself, $\hat{w}$ must belong to an ellipsoidal surface (e.g., a ball surface). Assume the surface can be written as $\{w|w^{\top}Rw=c^{2}\}$, then we can solve the maximizer $w^{*}=c\frac{R^{-1}Q^{-1}\hat{w}}{\sqrt{(Q^{-1}\hat{w})^{\top}R^{-1}(Q^{-1}\hat{w})}}$. To let $w^{*}=\hat{w}$, we need $Q^{-1}\hat{w}$ and $R\hat{w}$ have the same direction, which is equivalent to $\hat{w}$ is the eigenvector of $R^{-1}Q^{-1}$.
2. Proof of sufficiency: When $h$ is linear and $\mathcal{F}$ is also linear, the decision-maker can directly deploy $h=f$ to ensure the maximization of $\text{DW}(f),\text{SF}(f),\text{AW}(f)$. Next, according to the above derivation, $\text{IMP}(f)$ is also maximized.
Proof of Prop. 4.5
@@
1. Proof of the necessity: Since the maximization of the decision-maker welfare needs $f=h$, then $h\in\mathcal{F}$ is needed. Then the agents need to best respond to $f$ to ensure the maximization of social welfare. If there is an arbitrary agent who does not best respond to $f=h$, this means she misses the largest possible improvement and the overall maximization of social welfare is failed. Thus, according to Prop. 3.3, $h$ itself must be a polynomial of $X$ with at most order $K$ to guarantee agents to best respond. Finally, to ensure the maximization of $\text{IMP}(f)$, $\mathcal{F}$ must be linear.
@@
1. Proof of the necessity: If the condition does not hold this means there must exist a cost function $c(x,x^{\prime})$, under which we can find a region $A\subset\text{dom}(X)$ where $Q^{f^{*}}_{x}$ increases while $h$ decreases inside $A$. Thus, if we specify a cost function and an agent to satisfy the requirement that the agent can only change her features within $A$ and when $P_{X}$ is any continuous distribution, then $f^{*}$ is unsafe.
2. Proof of the sufficiency: We know for each dimension $i$, $Q^{f^{*}}_{x}$ improves at the same direction as $h$ and only the magnitude differs. This means agents will never be misguided by $Q^{f^{*}}_{x}$, thereby ensuring the safety. ∎
B.3 Proof of Thm. 4.7
@@
Appendix C Plots with error bars